### v1

# A Measurement of Solar Neutrino Flux using PandaX-4T Electron Recoil Data Xiaoying Lu

A Measurement of Solar Neutrino Flux using PandaX-4T Electron Recoil Data
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Abdusalam Abdukerim
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zihao Bo
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xun Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Chen Cheng
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Zhaokan Cheng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Xiangyi Cui
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Yingjie Fan
Department of Physics,Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Changbo Fu
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Mengting Fu
School of Physics, Peking University, Beijing 100871, China
Lisheng Geng
School of Physics, Beihang University, Beijing 102206, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
School of Physics and Microelectronics, Zhengzhou University, Zhengzhou, Henan 450001, China
Karl Giboni
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Linhui Gu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Chencheng Han
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Ke Han
ke.han@sjtu.edu.cnSchool of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Changda He
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Di Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yanlin Huang
School of Medical Instrument and Food Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhou Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Yu Hou
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Chenxiang Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Shuaijie Li
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Tao Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnTsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Wenbo Ma
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Xuyang Ning
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Zhicheng Qian
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Nasir Shaheed
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Xiaofeng Shang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Lin Si
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Andi Tan
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Yi Tao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xiuli Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Zhou Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Mengmeng Wu
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Weihao Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jingkai Xia
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Mengjiao Xiao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiang Xiao
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Pengwei Xie
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Binbin Yan
yanbinbin@sjtu.edu.cnTsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai, 519082, China
Jijun Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yong Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yukun Yao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Xinning Zeng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Dan Zhang
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Minzhen Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Shibo Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Zhang
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yuanyuan Zhang
Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Li Zhao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Qibin Zheng
School of Medical Instrument and Food Engineering, University of Shanghai for Science and Technology, Shanghai 200093, China
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Ning Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Xiaopeng Zhou
School of Physics, Beihang University, Beijing 102206, China
Yong Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Yubo Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
(January 13, 2024)
Abstract
The proton-proton () fusion chain dominates the neutrino production from the Sun. The uncertainty of the predicted neutrino flux is at the sub-percent level, whereas that of the best measurement is . Furthermore, no direct measurement has been made via electron scattering below 165 keV recoil energy. Using the PandaX-4T commissioning data with
0.63 tonneyear of exposure, we perform the first measurement of solar neutrinos in the recoil energy range from 24 to 144 keV. The neutrino flux is measured to be , consistent with Standard Solar Model and existing measurements.
According to the Standard Solar Model (SSM) [1, 2, 3], approximately 99% of the solar power comes from a series of reactions fusing hydrogen into helium.
Neutrinos emitted in the initial fusion of two protons to a deuteron constitute roughly 91% of the total solar neutrino flux. These neutrinos are commonly referred to as the neutrinos.
The sub-dominant neutrinos from electron capture of Be account for an additional 7% of the flux.
The weakly interacting neutrinos carry original information from solar fusion reactions.
The precise measurement of solar neutrino flux is essential for verifying the SSM and shedding light on the solar metallicity puzzle [4].
Solar neutrinos are also crucial for neutrino physics, especially for understanding the matter effect of neutrino oscillation [5].
The solar luminosity has been measured to a 0.6% precision.
In contrast, the neutrino flux has only been measured by Borexino with a precision of with recoil energy threshold at 165 keV [6, 7].
A number of multi-tonne-scale liquid xenon (LXe) experiments including PandaX-4T [8], LZ [9] and XENONnT [10] are under operation to search for dark matter particles, coherent scattering of solar neutrinos on xenon nuclei [11, 12], and possible abnormal magnetic moments of neutrinos [13, 14, 15] in the few or few-tens of keV-scale electron or nuclear recoil energy.
By design, these detectors also effectively cover over an electron recoil energy up to several hundred keV, spanning the majority of region of interest (ROI) for the neutrino flux.
Figure 1: Neutrino-electron elastic scattering spectrum of solar neutrinos with xenon electrons.
If the binding energy of atomic electrons (AE) is considered (red curve), the scattering event rate in our ROI is lower than the free electron (FE) scenario (green curve).
Adopted from Ref. [16].
The neutrino and electron interact via the electroweak force through the exchange of a Z or W boson, the latter of which is only possible for an electron neutrino.
The solar neutrinos have a continuous energy spectrum with an endpoint at 420 keV.
The Be neutrinos are mono-energetic at 384 keV (approximately 10%) and 862 keV (90%).
The expected event rate per unit recoil energy is
where is the number of target electrons, is the neutrino flux as a function of neutrino energy, () is the oscillation probabilities of into the flavor j, and is the differential cross-section.
Fig. 1 shows the energy spectrum of solar and Be neutrinos in a LXe detector, as given in Ref. [16].
When the xenon atomic effects are considered (adopted later in this paper), the rate in the ROI is suppressed by a few percent compared to the free electron scenario.
In this paper, we report the first measurement of the neutrino flux via their elastic scattering on electrons using the PandaX-4T detector located in the B2 hall of the China Jinping Underground Laboratory [17, 18].
The sensitive target of PandaX-4T has 3.69 tonne liquid xenon in a cylindrical dual-phase xenon time projection chamber (TPC) with 118.5 cm in diameter and 116.8 cm in height [19].
The prompt scintillation photons () and the delayed electroluminescence photons () are produced when a given energy is deposited in the sensitive volume.
Both and signals are recorded by the top and bottom photomultipliers (PMT) arrays, which have 169 and 199 Hamamatsu 3-inch PMTs installed.
Detailed discussions of the PandaX-4T detector can be found in Ref. [8].
Figure 2: The energy resolution using characteristic peaks at 41.6 keV, 163.9 keV, 208.1 keV, and 236.1 keV.
The red curve shows the fit function + 0.0061.
This analysis selects different data from the existing analyses of the first commissioning data release [8, 11, 20].
The energy ROI is from 24 keV to 144 keV, with more than 60% of the solar neutrinos included.
The lower boundary is chosen to be above the dark matter particle search region and the upper boundary to avoid the prominent 163.9 keV peak from Xe produced in the neutron calibration runs.
In the ROI, the detector noise has a negligible effect on both and signals.
Therefore, we can recover approximately 9.5% exposure previously excluded in low energy analysis due to elevated detector noise.
On the contrary, 8.4 days of data following a neutron calibration run is removed because of a high concentration of activated Xe and I.
In the end, a total of 86.5 days is used for this analysis.
The PMT gains are calibrated with a newly implemented “rolling gain” approach for PMTs, which calculates PMT gains by fitting single photoelectron (SPE) spectra in individual PMTs run by run, adopted already in recent Ref. [19].
Our previous analyses calculated PMT gains with weekly light-emitting-diode (LED) calibration.
To avoid biases in our data selections, cuts are first determined with LED-gain calibration data, validated with 16% of rolling gain physics data, and finally applied to the entire data set.
The quality cut variables are inherited from low energy analysis [8], but cut parameters are modified
to suit the different energy windows.
The main difference is the relaxation of the top and bottom PMT charge ratio requirement in the signal selection to accommodate the top saturation.
The total quality cut efficiency is .
The solar neutrino signals scattering on electrons are primarily single-site (SS) events within the ROI.
The identification of SS events follows the same procedure as in Ref. [8].
The SS efficiency is , calculated using the Rn calibration data.
A consistent efficiency is obtained using simulated data.
The horizontal position reconstruction follows the same procedure in Ref. [19], where de-saturated waveforms and improved optical Monte Carlo simulation are used.
The reconstruction uniformity is confirmed with the diffusive Kr calibration source injected into the TPC [21].
The mono-energetic peak of Kr is also used to generate an energy response map, and the corresponding correction procedure follows Ref. [8].
The energy in LXe TPC is reconstructed as in Refs. [22, 23].
The energy spectrum is further corrected using a quadratic function between reconstructed energy and true energy of characteristic peaks at 41.6 keV (), 163.9 keV (), 208.1 keV (), and 236.1 keV ( and ).
After the correction, the residual offset of energy peaks in the ROI is smaller than 1 keV, which is considered as the systematic uncertainty of the energy reconstruction.
The relative energy resolution at 24 keV (144 keV) is 8.8% (3.9%), as shown in Fig. 2.
Figure 3: The spatial distributions of the selected physics events in vs. (top) and vs. (bottom).
The dashed lines show the boundary of the fiducial volume, and blue (gray) dots represent events inside (outside) the fiducial volume.
The fiducial volume (FV) cuts and events in the ROI are shown in Fig. 3 as dashed lines and blue dots, respectively.
The FV selection is the same as in the low energy analysis geometrically, and a total of tonne of LXe in the center of the TPC is used.
The slight difference with respect to Ref. [8] is from the event position reconstruction.
The background sources in the ROI include radon and krypton impurities in liquid xenon, detector materials, and radioactive xenon isotopes.
The expected background sources are listed in Table 1 and described below.
Low energy background events from Rn chain are mainly the decay of Pb to the ground state of Bi [19].
The event ratio, defined as the fraction of Pb events in the ROI per decay, is determined from a dedicated calibration run with Pb events accumulated in about 11 days.
The event ratio is measured to be , with the uncertainty estimated based on the difference between our data and a recent theoretical prediction to accommodate the non-unique first-forbidden nature of the decay [24].
The number of Pb events in the ROI is .
The decay of Bi is a sub-dominate background component because it is rejected effectively by the Bi–Po coincidence analysis.
The Bi–Po pairs are identified by an event following a event in the subsequent 5 ms window.
In the time window, the number of randomly paired events is negligible.
We remove the identified Bi–Po pairs from our spectrum, and the contribution from Bi is no longer included in the spectrum fit.
Similar to Pb, the expected contribution from decay of Pb in Rn decay chain is determined from spectrum fit from higher energy analysis [19].
The maximum deviation of Pb/Po throughout the calibration period and stable decay period is considered as the systematic uncertainty.
The expected number of events in the ROI is .
The concentration of Kr is 0.52 0.27 parts per trillion [19].
The expected contribution of Kr is 489 254 in the ROI.
The levels of radioactivity from PMTs and detector vessels have been determined from the wide energy spectrum fit in Ref. [20].
The expected contribution in ROI is events.
The expected background from Xe is in the ROI, given by the half-life measured by PandaX-4T experiment with an exposure of 15.5 kgyear of Xe isotope [20].
The contributions of Xe, Xe, I, and Xe are free and fitted in the final spectral analysis, as there are no constraints in our data outside the ROI.
Xe from neutron activation during calibration contributes to the ROI with a distinctive rising slope starting at 81 keV in the energy spectrum.
Cosmogenic Xe, introduced when some xenon exposed at the Earth’s surface was filled into the detector during the data taking, contributes to the background with K-shell electron capture around 33.2 keV.
Neutron-activated Xe decays quickly to
the relatively long-lived I with a half-life of 59.4 days.
I electron capture results in energy release of 67.3 keV (80% K-shell), 40.4 keV (16% L-shell), and 36.5 keV (3.5% M-shell) in the ROI.
Thanks to the efficient removal capability of the xenon purification and circulation system, the effective lifetime of I in the TPC is significantly reduced and cannot be constrained by the isotopic lifetime.
Two-neutrino double-electron capture of Xe (abundance = 0.095%, Q = 2857 keV) deposits energy within our ROI at 64.3 keV and 32.3 - 37.3 keV [25, 26].
Table 1: Expected and fitted background and signal counts in the ROI.
Components
Expectedcounts
Fittedcounts
Pb
1865 110
1845 113
Pb
276 71
270 80
Kr
489 254
405 249
Material
683 27
681 27
Xe
1009 46
999 47
Xe
free
4751 136
Peak 1: [32-41] keV
free
119 27
Peak 2: [64-68] keV
free
268 37
+Be neutrino
297 260
The spectrum fit is based on a binned likelihood procedure using the RooFit package [27].
Each background component is simulated using BambooMC [28], a Monte Carlo simulation package based on the Geant4 framework [29].
The contributions from Pb, Pb, Kr, detector material, and Xe are constrained using the uncertainties of the expected counts as described earlier and listed in Table 1.
Multiple mono-energetic peaks around 67 keV (33 keV) are grouped and modeled using a Gaussian function.
The means of the Gaussian functions are free in the fit, while the sigmas are constrained by a fractional uncertainty of 20% of the measured resolutions in Fig. 2.
The contributions of Xe and two Gaussian peaks are free-floating in the fit.
The sum of solar and Be neutrino signals are combined into a single fit component.
The spectrum fit result is shown in Fig. 4.
The number of solar +Be neutrino signals is 297 260.
Consistent results are obtained with a binned likelihood fit similar to that implemented in Ref. [20].
The fitted results of each background are listed in Table 1.
The background contributions from the material, Kr, Pb, and Pb are consistent with the expected values.
The fitted uncertainty includes a statistical () and a systematic () uncertainty of 107 and 237 events, respectively, extracted by comparing the signal uncertainties with and without the penalty terms for the systematic constraints.
A similar ad hoc approach is used to extract the contribution from individual systematic effects later, assuming that extracted uncertainty is uncorrelated with the rest.
Xe (free) is the largest background source, however its contribution to is relatively insignificant at 43 events due to its characteristic spectral shape.
Table 2: Summary of contributions to systematic uncertainties. The contribution of each constrained parameter is extracted by turning on/off the corresponding penalty term in the fitter, assuming no correlation with other constraints. As a result, the quadrature combination of those rows is slightly different from the subtotal obtained directly by the fitter with all correlations properly built in.
Components
Counts
Kr
203
Pb
87
Pb
69
Material
21
Xe
19
Peak 1 width
37
Peak 2 width
33
Data selection
30
Subtotal
237
Energy scale
111
Energy resolution
19
Fit range
32
Pb spectrum
82
Pb spectrum
17
Kr spectrum
6
Subtotal
144
Total
277
Table 2 lists all the components of the systematic uncertainties.
Among the constrained backgrounds, Kr, Pb and Pb dominate the contribution to with an uncertainty of 203, 87, and 69 events, respectively.
The constraints of two Gaussian peak widths give an uncertainty of 37 and 33 events, respectively.
The uncertainties of data selection, including FV, quality cuts, and SS fraction, affect the uncertainty of the signal number proportionally with 30 events.
Additional sources of systematic uncertainties are evaluated manually ().
The uncertainty from the reconstructed energy scale is introduced by shifting the energy spectrum by keV and comparing the new fit results with the baseline fit, with a result of 111 events. To evaluate the contribution due to the energy resolution uncertainty,
we perform the fit using PDFs with different energy resolution values.
A difference in energy resolution, where is from the aforementioned resolution function, propagates into an uncertainty of 19 events.
The fit range uncertainty is evaluated to be 32 events by varying the ROI to 25 – 143 keV and 23 – 145 keV.
The uncertainty introduced by Pb spectrum shape is calculated by changing the theoretical spectrum in the baseline fit to the shape measured in calibration data, leading to a result of 82 events.
Similarly, the spectral shape differences between Geant4 [28] and a recent theoretical calculation of Pb and Kr [24] are also considered, and the introduced systematic uncertainties are found to be 17 and 6 respectively.
In total, we get an overall systematic uncertainty of 277 events with 144 from .
Figure 4: (Color online) Result of the spectrum fit and the corresponding residual plot in the bottom panel. The solar +Be neutrino is shown as the red line.
The constrained and free-floating backgrounds are shown in solid and dashed lines, respectively.
The (number of degrees of freedom, defined as the number of bins minus the number of free fit parameters) is shown in the legend.
The final number of +Be neutrinos is from our analysis.
The corresponding solar neutrino flux is , taking into account of the expected ratio of the and Be fluxes from SSM and their contributions in the ROI.
The result is consistent with the SSM expectation [4] and the Borexino measurement [7], as shown in Fig. 5.
This analysis represents the first measurement for solar neutrino at low recoil energy and demonstrates the potential of multi-tonne-scale LXe detectors for solar neutrino studies in a completely new energy window.
With an optimized online cryogenic distillation system, we expect to reduce the radon and krypton concentration by a factor of 1.8 or more [30].
Additional efforts, such as replacing TPC materials and circulation pumps, are being implemented for background control.
Additional Rn calibration data and improved detector response with upgraded PMT readout circuit boards are expected to significantly reduce systematic uncertainties.
Assuming radon and krypton levels of 3.5 Bq/kg and 0.25 ppt, respectively, negligible activation of xenon isotopes, and imposing a 5% constraint on all backgrounds, PandaX-4T can measure the solar neutrinos with an uncertainty of 30% with 6 tonneyear exposure.
Figure 5: The red dot and line show the result of solar neutrino flux, compared with the measurement from Borexino [7] and PandaX-4T projected measurement with 6 tonneyear exposure.
The vertical dashed line is the average neutrino flux from the high-metallicity and low-metallicity SSM [4].
This project is supported in part by the grants from National Science Foundation of China (Nos. 12090060, 12090063, 12105052, 12005131, 11905128, 11925502), and by the Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100).
We thank the support from the Double First Class Plan of Shanghai Jiao Tong University.
We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, Tencent, and New Cornerstone Science Foundation in China. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Vinyoles et al. [2017]N. Vinyoles, A. M. Serenelli, F. L. Villante, S. Basu,
J. Bergström, M. C. Gonzalez-Garcia, M. Maltoni, C. Peña Garay, and N. Song, A new Generation of Standard Solar Models, Astrophys. J. 835, 202 (2017), arXiv:1611.09867 [astro-ph.SR]
de Holanda and Smirnov [2003]P. C. de Holanda and A. Y. Smirnov, LMA MSW solution of the
solar neutrino problem and first KamLAND results, JCAP 02, 001, arXiv:hep-ph/0212270 .
Bellini et al. [2014]G. Bellini et al. (BOREXINO), Neutrinos from the primary
proton–proton fusion process in the Sun, Nature 512, 383 (2014).
Zhou et al. [2021]X. Zhou et al. (PandaX-II), A Search for Solar Axions and Anomalous Neutrino
Magnetic Moment with the Complete PandaX-II Data, Chin. Phys. Lett. 38, 011301 (2021), [Erratum: Chin.Phys.Lett. 38, 109902 (2021)], arXiv:2008.06485 [hep-ex] .
Billard et al. [2014]J. Billard, L. Strigari, and E. Figueroa-Feliciano, Implication
of neutrino backgrounds on the reach of next generation dark matter direct
detection experiments, Phys. Rev. D 89, 023524 (2014), arXiv:1307.5458 [hep-ph]
Yan et al. [2023]X. Yan et al. (PandaX), Searching for Two-Neutrino and Neutrinoless Double Beta
Decay of Xe with the PandaX-4T Experiment (2023), arXiv:2312.15632
[nucl-ex] .
Szydagis et al. [2011]M. Szydagis, N. Barry,
K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, NEST: A
Comprehensive Model for Scintillation Yield in Liquid Xenon, JINST 6, P10002, arXiv:1106.1613 [physics.ins-det] .
Haselschwardt et al. [2020]S. J. Haselschwardt, J. Kostensalo, X. Mougeot, and J. Suhonen, Improved calculations of
beta decay backgrounds to new physics in liquid xenon detectors, Phys. Rev. C 102, 065501 (2020), arXiv:2007.13686 [hep-ex] .