### v2

# Measurement of Solar Neutrino Flux using Electron Recoil Data from PandaX-4T Commissioning Run Xiaoying Lu

Measurement of Solar Neutrino Flux using Electron Recoil Data from PandaX-4T Commissioning Run
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Abdusalam Abdukerim
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zihao Bo
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xun Chen
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chen Cheng
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Zhaokan Cheng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Xiangyi Cui
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Yingjie Fan
Department of Physics,Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Lisheng Geng
School of Physics, Beihang University, Beijing 102206, China
Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chencheng Han
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Ke Han
ke.han@sjtu.edu.cnSchool of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Changda He
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Di Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhou Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yu Hou
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chenxiang Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shuaijie Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Tao Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Wenbo Ma
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xuyang Ning
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhicheng Qian
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Nasir Shaheed
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Xiaofeng Shang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Lin Si
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yi Tao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xiuli Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xu Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Zhou Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Mengmeng Wu
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Weihao Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuan Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Mengjiao Xiao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiang Xiao
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Binbin Yan
yanbinbin@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai, 519082, China
Yong Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Youhui Yun
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xinning Zeng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Minzhen Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shibo Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wei Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yuanyuan Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
Li Zhao
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Ning Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai, 200240, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xiaopeng Zhou
School of Physics, Beihang University, Beijing 102206, China
Yubo Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhizhen Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
(June 14, 2024)
Abstract
The proton-proton () fusion chain dominates the neutrino production from the Sun.
The uncertainty of the predicted neutrino flux is at the sub-percent level, whereas that of the best measurement is .
In this paper, we present the first result to measure the solar neutrinos in the electron recoil energy range from 24 to 144 keV, using the PandaX-4T commissioning data with 0.63 tonneyear exposure.
The neutrino flux is determined to be , consistent with Standard Solar Model and existing measurements, corresponding to a flux upper limit of at 90% C.L..
According to the Standard Solar Model (SSM) [1, 2, 3], approximately 99% of solar power comes from a series of reactions fusing hydrogen into helium.
Neutrinos emitted in the initial fusion of two protons to a deuteron constitute roughly 91% of the total solar neutrino flux. These neutrinos are commonly referred to as the neutrinos, with a theoretical flux uncertainty at the sub-percent level.
The sub-dominant neutrinos from electron capture of 7Be account for an additional 7% of the flux.
The precise measurement of solar neutrino flux is essential for verifying the SSM and shedding light on the solar metallicity puzzle [4].
Solar neutrinos are also crucial for neutrino physics, especially for understanding the matter effect of neutrino oscillation [5].
The first observation of neutrino flux is achieved using radiochemical material of gallium by GALLEX/GNO [6] and SAGE [7]. The first real-time detection is made by Borexino with now a state-of-the-art precision of and an electron recoil energy threshold of 165 keV [8, 9].
In this paper, we present a new result to measure neutrino flux in real-time by PandaX-4T within a previously unaccessible recoil energy window between 24 to 144 keV. It also represents the first such measurement using a detector primarily designed for dark matter direct detection.
A number of multi-tonne-scale liquid xenon (LXe) experiments, including PandaX-4T [10], LZ [11], and XENONnT [12] are under operation to search for dark matter particles, coherent scattering of solar neutrinos on xenon nuclei [13, 14], and possible abnormal magnetic moments of neutrinos [15, 16, 17] in the few or few-tens of keV-scale electron or nuclear recoil energy.
By design, these detectors also effectively cover an electron recoil energy up to several hundred keV, spanning over most of the energy region of the neutrino.
Figure 1: Neutrino-electron elastic scattering spectrum of solar neutrinos with liquid xenon.
If the binding energy of atomic electrons (AE) is considered [18] (red curve), the scattering event rate in our region of interest is lower than the free electron (FE) scenario (green curve). The more complete treatment using the relativistic random phase approximation (RRPA) is shown in cyan, but only for a recoil energy less than 30 keV.
The solar neutrinos have a continuous energy spectrum with an endpoint at 420 keV.
The 7Be neutrinos are mono-energetic at 384 keV (approximately 10%) and 862 keV (90%).
The neutrino and electron interact via the electroweak force through the exchange of a Z or W boson, the latter of which is only possible for an electron neutrino.
The expected electron recoil event rate per unit recoil energy is
where is the number of target electrons, is the neutrino flux as a function of neutrino energy, () is the oscillation probabilities of into flavor j and is the differential cross-section.
Fig. 1 shows the electron recoil energy spectrum of solar and 7Be neutrinos in a LXe detector, as given in Ref. [18].
When the xenon atomic effects are considered (adopted in this paper), the rate in the region of interest (ROI) is suppressed by a few percent compared to the free electron scenario.
The PandaX-4T detector is located in the B2 hall of the China Jinping Underground Laboratory [19, 20].
The sensitive target of PandaX-4T contains 3.69 tonne of liquid xenon in a cylindrical dual-phase xenon time projection chamber (TPC) with 118.5 cm in diameter and 116.8 cm in height [21].
The prompt scintillation photons () and the delayed electroluminescence photons () are produced when a given energy is deposited in the sensitive volume.
Both and signals are recorded by the top and bottom photomultipliers (PMT) arrays, which have 169 and 199 Hamamatsu 3-inch PMTs installed.
Detailed discussions of the PandaX-4T detector can be found in Ref. [10].
Figure 2: The energy resolution using characteristic peaks at 41.6 keV, 163.9 keV, 208.1 keV, and 236.1 keV.
The red curve shows the fit function + 0.0061.
This analysis selects different data samples compared to existing analyses of the commissioning data release [10, 13, 22].
The energy ROI is from 24 keV to 144 keV, with more than 60% of the solar neutrino events included.
The lower boundary is chosen to be above the dark matter particle search region and the upper boundary is to avoid the prominent 163.9 keV peak from 131mXe produced in the neutron calibration runs.
In the ROI, the detector noise has a negligible effect on both and signals.
Therefore, we can recover approximately 9.5% of exposure previously excluded in the dark matter analysis due to elevated detector noise [10].
On the contrary, 8.4 days of data following a neutron calibration run is removed because of a high concentration of activated 133Xe and 125I.
In the end, a total data set of 86.5 days is used for this analysis.
The PMT gains are calibrated with a newly implemented “rolling gain” approach, which fits single photoelectron (SPE) spectra in individual PMTs run by run, adopted already in recent Ref. [21].
Our previous analyses calculated PMT gains with weekly light-emitting-diode (LED) calibration.
To avoid biases in our data selections, selection cuts are first determined with LED-gain calibration data, validated with about 13.5 days of rolling gain physics data, and finally applied to the entire data set.
The quality cut variables are inherited from the dark matter analysis [10], but cut parameters are modified
to suit the differing energy window.
The main difference is the relaxation of the top and bottom PMT charge ratio requirement in the signal selection to accommodate the top saturation.
The total quality cut efficiency is .
The scattering of solar neutrinos on electrons is primarily single-site (SS).
The identification of SS events follows the same procedure as in Ref. [10].
The SS efficiency is , calculated using the 220Rn calibration and consistent with simulation.
The horizontal position reconstruction follows the same procedure in Ref. [21], where de-saturated waveforms and improved optical Monte Carlo simulation are used.
The reconstruction uniformity is confirmed with the diffusive 83mKr calibration source injected into the TPC [23].
The mono-energetic peak of 83mKr is also used to generate an energy response map. The corresponding correction procedure follows Ref. [10].
The energy in LXe TPC is reconstructed as in Refs. [24, 25].
The energy spectrum is further corrected using a quadratic function between reconstructed energy and true energy of characteristic peaks at 41.6 keV (), 163.9 keV (), 208.1 keV (), and 236.1 keV ( and ).
After the correction, the residual offset of energy peaks in the ROI is smaller than 1 keV, which is considered as the systematic uncertainty of the energy reconstruction.
The relative energy resolution at 24 keV (144 keV) is 8.8% (3.9%), as shown in Fig. 2.
Figure 3: The spatial distributions of the selected physics events in vs. (top) and vs. (bottom).
The dashed lines show the boundary of the fiducial volume, and blue (gray) dots represent events inside (outside) the fiducial volume.
The fiducial volume (FV) cuts and events in the ROI are shown in Fig. 3 as dashed lines and blue dots, respectively.
The FV selection is the same as in the dark matter analysis geometrically, and a total of tonne of LXe in the center of the TPC is used.
The slight difference with respect to Ref. [10] is from the event position reconstruction.
The primary challenge in this analysis lies in the accurate rate estimation of various background components with smooth energy spectra that resemble the shape of the solar signals.
The background sources in the ROI include radon and krypton impurities in liquid xenon, detector materials, and radioactive xenon isotopes.
The expected background sources are listed in Table 1 and described below.
One major background in the ROI arises from the progeny of the internal 222Rn – the decay of 214Pb to the ground state of 214Bi [21].
The ratio of the events in the ROI to the full SS energy spectrum of 214Pb decay is determined by a dedicated 222Rn injection calibration run, with 214Pb SS events accumulated over approximately 11 days.
The event ratio is measured to be , with the uncertainty estimated based on the difference between our data and a recent theoretical prediction to accommodate the non-unique first-forbidden nature of the decay [26].
The corresponding number of 214Pb events in the ROI is , in which the uncertainty combines both the fitted uncertainty of the overall 214Pb decay rate in the physics data [21] and the uncertainty of the event ratio within our ROI. Also along the decay chain of 222Rn, the decay of 214Bi can be identified effectively using the 214Bi–214Po coincidence, as the half-life of 214Po is 164 s.
In our analysis, when a 214Bi-like event is found, we reject the event if an event is found within the following 5 ms window. The residual 214Bi is no longer included in the spectrum fit. The loss of live time due to random coincidence is negligible.
The expected contribution from the decay of 212Pb in the 220Rn decay chain follows the approach in Ref. [21].
The ratio between 212Pb and the subsequent 212Po events is determined using a 220Rn injection calibration run. Then using the 212Po events identified in the physics data,
212Pb events in the ROI is estimated to be , where the uncertainty is dominated by the variation of 212Pb/212Po ratio during the 220Rn injection run.
The concentration of 85Kr is estimated to be 0.52 0.27 parts per trillion [21], leading to 489 254 events in the ROI.
The levels of radioactivity from PMTs and detector vessels have been determined from the wide energy spectrum fit in Ref. [22], resulting in events. The expected background from 136Xe is in the ROI, given by the half-life measured by the PandaX-4T experiment with an exposure of 15.5 kgyear of 136Xe isotope [22].
The contributions of 133Xe, 127Xe, 125I, and 124Xe are free and fitted in the final spectral analysis, as they cannot be constrained by our data outside the ROI.
The energy spectrum of 133Xe, activated during a neutron calibration run, has a distinctive rising slope starting at 81 keV.
Cosmogenic 127Xe, introduced when some xenon exposed at the Earth’s surface was filled into the detector, contributes to the background with K-shell electron capture around 33.2 keV.
Neutron-activated 125Xe decays quickly to
the relatively long-lived 125I with a half-life of 59.4 days.
125I electron capture can release a total energy of 67.3 keV (K-shell), 40.4 keV (L-shell), and 36.5 keV (M-shell) [27]. The reduction rate of 125I observed in the TPC is significantly more rapid than the natural lifetime of 125I, likely due to the removal by the xenon circulation and purification system.
Two-neutrino double-electron capture of 124Xe (abundance = 0.095%, Q = 2857 keV) deposits energy within our ROI at 64.3 keV (KK-shell) and 32.3 - 37.3 keV (KL, KM, and KN-shells) [28, 27]. Therefore, 125I (124Xe) events in the data are modeled as three (four) Gaussian peaks with relative areas fixed by the capture fractions and widths fixed by the resolution function (Fig. 2), but the overall normalization as a free fit parameter.
Table 1: Expected and fitted background and signal events in the ROI. Note that the fitted uncertainties contains both the statistical and some systematic components. See text for details.
Components
Expectedevents
Fittedevents
214Pb
1865 110
1849 113
212Pb
276 71
271 80
85Kr
489 254
423 249
Materials
683 27
682 27
136Xe
1009 46
1002 47
133Xe
free
4767 135
124Xe
free
317 63
125I
free
31 57
127Xe
free
59 23
+7Be neutrino
231 257
The spectrum fit is based on a binned likelihood procedure using the RooFit package [29].
Each background component is simulated using BambooMC [30], a Monte Carlo simulation package based on the Geant4 framework [31].
The contributions from 214Pb, 212Pb, 85Kr, detector materials, and 136Xe are constrained using the uncertainties described earlier and listed in Table 1.
The normalizations of 133Xe, 127Xe, 125I, and 124Xe are kept free in the fit.
The sum of solar and 7Be neutrino signals are combined into a single fit component.
The result of the spectrum fit is shown in Fig. 4.
The number of solar + 7Be neutrino signals is 231 257 events.
Consistent results are obtained with a binned likelihood fit similar to that implemented in Ref. [22].
The fitted results of each background are listed in Table 1.
The best-fit background contributions from detector materials, 85Kr, 214Pb, and 212Pb are very close to the input nominal values, indicating that the spectrum fit does not try to drive these background levels up and down, since their shape characters are similar.
The total uncertainty of the + 7Be neutrino signals is broken down into three terms (Table 2), the statistical uncertainty (), the systematic uncertainty incorporated in the likelihood fit with nuisance parameters (), and additional systematic uncertainty evaluated by hand (). We shall discuss them in turn below.
The statistical uncertainty is determined to be 113 events, by re-fitting the data with all constrained parameters fixed to their best-fit values without the penalty terms.
The overall size of is then evaluated as to be 231 events, representing the residual component of the fit uncertainty. To quantify the contribution to due to a given component (assuming that it is uncorrelated with the rest), we use a similar approach by refitting the data by only “tuning on” the nuisance parameter of this component, and all other nuisance parameters fixed to their best-fit values. The resulting fit uncertainty is then subtracted by in quadrature. Individual contributions thus evaluated are also summarized in Table 2. Among all constrained backgrounds, 85Kr, 214Pb and 212Pb dominate the contribution to with an uncertainty of 202, 87, and 69 events, respectively.
The uncertainties of data selection, including the FV, quality cuts, and the SS fraction, affect the signal number proportionally with an estimated uncertainty of 29 events.
Additional systematic uncertainties are evaluated manually.
The uncertainty from the reconstructed energy scale is determined by comparing the baseline fit result with new fit results, obtained by shifting the data energy spectrum by keV.
The larger difference led by the shifts is 142, which is used as the uncertainty.
To evaluate the contribution due to the energy resolution uncertainty, we perform the fit by imposing different background spectra with different energy resolution.
A difference in the energy resolution function propagates into an uncertainty of 19 events.
The fit range uncertainty is evaluated to be 29 events by varying the ROI to 25 – 143 keV and 23 – 145 keV.
The uncertainty introduced by the 214Pb spectrum shape estimated to be 84 events, by comparing the theoretical spectrum in the baseline fit to the shape directly measured in the calibration data.
Similarly, the spectral shape differences between the Geant4 [30] prediction and a recent theoretical calculation of 212Pb and 85Kr [26] are also considered, and the systematic uncertainties are found to be 18 and 5, respectively.
We also fit the data by taking the half-life of 136Xe from EXO-200 [32], the most precise result so far, as the input value.
This results in a difference of 16 events in comparison to our baseline fit, which is also taken as a component of systematic uncertainty. In total, is 170 events.
Table 2: Summary of contributions to the + 7Be neutrino signal uncertainties. The contribution of each constrained parameter is extracted by turning on/off the corresponding penalty term in the fitter, assuming no correlation with other constraints. As a result, the quadrature combination of those rows is slightly different from the subtotal obtained directly by the fitter with all correlations properly built in.
Components
Counts
113
85Kr
202
214Pb
87
212Pb
69
Material
21
136Xe
19
Data selection
29
Subtotal
231
Energy scale
142
Energy resolution
19
Fit range
29
214Pb spectrum
84
212Pb spectrum
18
85Kr spectrum
5
136Xe half-life
16
Subtotal
170
Total
287
Figure 4: (Color online) Result of the spectrum fit and the corresponding residual plot in the bottom panel. The solar +7Be neutrino is shown as the red line.
The constrained and free-floating backgrounds are shown in solid and dashed lines, respectively.
Therefore, the final detected number of + 7Be neutrinos is events from our analysis.
The corresponding solar neutrino flux is , using the expected ratio of the to 7Be fluxes from SSM and their contributions to the ROI.
The result is consistent with the SSM expectation [4] and the Borexino measurement [9], as shown in Fig. 5.
Based on this, we obtain a neutrino flux upper limit of at 90% C.L..
This analysis represents the first result to directly measure the solar neutrinos at an electron recoil energy below 150 keV. It demonstrates the potential of multi-tonne-scale LXe detectors for solar neutrino studies in a completely new energy window.
With an optimized online cryogenic distillation system, we expect to reduce the radon and krypton concentration by a factor of 1.8 or more [33].
Additional efforts, such as replacing TPC materials and circulation pumps, are being implemented for background reduction.
Additional 222Rn calibration data and improved detector response with upgraded PMT readout circuit boards are expected to significantly reduce systematic uncertainties.
Assuming a 222Rn level of 3.5 Bq/kg, a krypton-to-xenon mole concentration of 0.25 ppt, and no short-lived xenon isotopes being activated, and imposing a 5% constraint on all backgrounds, PandaX-4T can measure the solar neutrinos with an uncertainty of 30% with 6 tonneyear exposure.
Figure 5: The solar neutrino flux measured by PandaX-4T (red), in comparison to those from gallium experiments [34] and Borexino [9].
The vertical dashed line is the average neutrino flux from the high-metallicity and low-metallicity SSM [4].
The PandaX-4T projection with 6 tonneyear exposure is indicated at the bottom (centered at the SSM).
This project is supported in part by the grants from National Science Foundation of China (Nos. 12090060, 12090063, 12105052, 12005131, 11905128, 11925502), the Office of Science and Technology, Shanghai Municipal Government (grant No. 22JC1410100),
and National Science Foundation of Sichuan Province (No. 2024NSFSC1371).
We thank the support from the Double First Class Plan of Shanghai Jiao Tong University.
We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Hongwen Foundation in Hong Kong, Tencent, and New Cornerstone Science Foundation in China. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Vinyoles et al. [2017]N. Vinyoles, A. M. Serenelli, F. L. Villante, S. Basu,
J. Bergström, M. C. Gonzalez-Garcia, M. Maltoni, C. Peña Garay, and N. Song, A new Generation of Standard Solar Models, Astrophys. J. 835, 202 (2017), arXiv:1611.09867 [astro-ph.SR]
de Holanda and Smirnov [2003]P. C. de Holanda and A. Y. Smirnov, LMA MSW solution of the
solar neutrino problem and first KamLAND results, JCAP 02, 001, arXiv:hep-ph/0212270 .
Anselmann et al. [1992]P. Anselmann et al. (GALLEX), Solar neutrinos observed by GALLEX at Gran
Sasso., Phys. Lett. B 285, 376 (1992).
Zhou et al. [2021]X. Zhou et al. (PandaX-II), A Search for Solar Axions and Anomalous Neutrino
Magnetic Moment with the Complete PandaX-II Data, Chin. Phys. Lett. 38, 011301 (2021), [Erratum: Chin.Phys.Lett. 38, 109902 (2021)], arXiv:2008.06485 [hep-ex] .
Billard et al. [2014]J. Billard, L. Strigari, and E. Figueroa-Feliciano, Implication
of neutrino backgrounds on the reach of next generation dark matter direct
detection experiments, Phys. Rev. D 89, 023524 (2014), arXiv:1307.5458 [hep-ph]
Szydagis et al. [2011]M. Szydagis, N. Barry,
K. Kazkaz, J. Mock, D. Stolp, M. Sweany, M. Tripathi, S. Uvarov, N. Walsh, and M. Woods, NEST: A
Comprehensive Model for Scintillation Yield in Liquid Xenon, JINST 6, P10002, arXiv:1106.1613 [physics.ins-det] .
Haselschwardt et al. [2020]S. J. Haselschwardt, J. Kostensalo, X. Mougeot, and J. Suhonen, Improved calculations of
beta decay backgrounds to new physics in liquid xenon detectors, Phys. Rev. C 102, 065501 (2020), arXiv:2007.13686 [hep-ex] .
Abdurashitov et al. [2009]J. N. Abdurashitov et al. (SAGE), Measurement of the solar neutrino capture rate
with gallium metal. III: Results for the 2002–2007 data-taking period, Phys. Rev. C 80, 015807 (2009), arXiv:0901.2200 [nucl-ex] .